im really not sure how to do this i have to find the domain and range of y=square root of 3+x. how do i find the domain and range? thanks
Answer by stanbon(57967)

find the domain and range of y=square root of 3+x. how do i find the domain and range?
---------
To find the domain you have to look for Real Number Values that x CANNOT take.
In your problem you have y=sqrt(3+x)
--------
That 3+x cannot be negative;it must be >=0
So solve 3+x>=0
x>-3
The domain is all Real Numbers >=3
-------------------------
For the range you need to determine the Real Number values y CAN take.
Since you have y=sqrt(3+x), y must be >=0
The range is All Real Numbers >=0
=============
Cheers,
Stan H.